OpenStudy (anonymous):

what is the steps to 7/8 divided by 3/5

OpenStudy (anonymous):

Write the first fraction down and the second one flipped upside down \[\frac{ 7 }{ 8 } \times \frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

Now multiply the top numbers together and bottom number together \[\frac{ 7 }{ 8 } \times \frac{ 5 }{ 3 }\]\[= \frac{ 35 }{ 24 }\]

OpenStudy (anonymous):

now simplify if possible, but its not here